# An approach based on metaheuristic algorithms to the timetabling problem in deregulated railway markets

### v1

@@ 3.3 General Scheme of the Optimization Process
Figure 4: General optimization scheme
@@ 1. Generating schedule requests:
Let $SS=\{S_{1},\dots,S_{|SS|}\}$ be the set of schedules, where each $S_{i}$ is a schedule. Each service in $S_{i}$ is composed of a set of stops. For each stop, Robin indicates the station ($Sta_{i}^{j}$), departure ($dt_{i}^{j}$) and arrival time ($at_{i}^{j}$). Thus, $S_{i}$ can be expressed as:
$S_{i}=\{(Sta_{i}^{1},dt_{i}^{1},at_{i}^{1}),\dots,(Sta_{i}^{|S_{i}|},dt_{i}^{|S_{i}|},at_{i}^{|S_{i}|})\}$ (8)
@@
2. Computing Departure time vector: Then, a one-dimension vector containing the departure times of the services in a sorted way is generated using $SS$. This vector is called $DT$ and can be expressed as:
@@
As mentioned above, with the original requested times, it was not feasible to schedule all three services. However, based on the new $ODT$ times this would be feasible, if a five minute minimum safety gap for each service is included. In order to reach this solution, $S_{1}$ has been advanced by half an hour and departure of service 2 has been delayed, also by thirty minutes.
It is important to remark that this might not be the optimal solution: for example, if RU fleeting service one is much more sensitive to penalties compared to the others, it would be better to delay service two and service three while not modifying the requested times of service 1.
@@
In the following example, based on the previous $ODT$ vector, a definitive schedule is devised. Even with the optimized departure times vector, it is possible to find conflicts between services, making it infeasible to schedule all the requests. For this reason, the heuristic for service planning is also applied. The Marey chart with the final schedule based on the $ODT$ previously shown for the example schedule can be seen in Figure 5. The modified schedule resulting from the $ODT$ vector is set out in Table 3.
@@
Table 3: Final timetable of rail services organized by station
Figure 5: Example schedule based on $ODT$
@@ 3.3.1 A Heuristic for Railway Service Scheduling
To address the exponential growth caused by checking all possible combinations, the following heuristic technique has been implemented to obtain the vector of trains to be scheduled. The operation of this heuristic is detailed in Algorithm 1. With the proposed method, services without conflicts are scheduled first (Line 5). On the other hand, for services with conflicts, an iterative process is carried out (Lines 7–12). In this phase, the service that provides the highest benefit is scheduled sequentially (Lines 8–9), and the possibility of scheduling any service that presents conflicts with the one selected is removed for future iterations (Lines 10–11).
@@
1:  Input: Proposed time vector
2:  Output: Scheduled services vector
3:  Initialize schedule with the times from the vector
4:  Determine conflicts between services
5:  Schedule the services without conflicts
6:  Calculate revenues associated with services that have conflicts
7:  while There are services with conflicts do
8:     Select the service with the highest associated benefit
9:     Schedule the selected service
10:     Identify services that conflict with the one scheduled
11:     Update the list of conflicting services by removing the ones now scheduled and their conflicts
12:  end while
@@ 4 Experiments and results
This section presents the results of metaheuristic algorithms for solving the timetabling problem in a liberalized railway market, first without hyperparameter optimization (Section 4.1), and then after optimization (Section 4.2) This enables evaluation of their effectiveness and configuration recommendations. The Python library used is MEALPY [9], as it is robust and efficient and implements a large number of cutting-edge population-based metaheuristic algorithms.
@@ 4.1 Tests without hyperparameter tuning
With the first experiment, all the selected metaheuristics were tested to maximize the IM revenue obtained based on the same 25 schedule requests. In this first test, no hyperparameter optimization was applied, selecting a standard 100-epoch optimization process and 20 individuals (for the population-based metaheuristics) and $100\cdot 20=2000$ epoch for the single-solution-based algorithms, i.e. SA. The default parameters for each algorithm are shown in Table 4. There are no parameters that can be adjusted in the implemented MEALPY version of Covariance Matriz Adaptation (CMA-ES), Grey Wolf Optimize (GWO), Whale Optimization Algorithm (WOA) and Hybrid Grey Wolf - Whale Optimization Algorithm (GWO-WOA). Thus, they are not mentioned in Table 4.
Algorithm Parameter Default Value Genetic Algorithm (GA) Crossover probability ($pc$) 0.95 Mutation probability ($pm$) 0.025 Selection method Tournament Particle Swarm Optimization (PSO) Local coefficient ($c1$) 2.05 Global coefficient ($c2$) 2.05 Positive constant (alpha) 0.4 Simulated Annealing (SA) Initial temperature 100 Cooling rate 0.99 Differential Evolution (DE) Weighting factor ($wf$) 0.1 Crossover rate ($cr$) 0.9 Ant Colony Optimization Continuous (ACOR) Sample count 25 Intensification factor 0.5 Zeta 1.0 Artificial Bee Colony (ABC) Food source limit 25
Table 4: Adjustable parameters in MEALPY for each algorithm
@@
Figure 7: Convergence curves
Figure 8: Convergence curves of Top 3 algorithms
@@
Figure 9: Scattered Boxplot of Population
Table 5 shows the results achieved for each algorithm for each one of the five different runs, ranked by the obtained revenue (higher to lower). The results from the table highlight the clear dominance of GA across all metrics. It consistently achieves the highest revenue over the majority of runs, demonstrating its effectiveness in optimizing the problem at hand. However, an interesting observation arises in runs three and five of GA, where scheduling more trains (nineteen in run five versus eighteen in run three) does not necessarily translate into higher revenue. This indicates that the quality of the solution is not solely dependent on the number of scheduled trains, but also on other factors such as minimizing schedule changes. This happens because, in order to schedule a higher number of services, it is often required to make more significant changes to the requested times, thereby incurring higher penalties. The metrics $\Delta DT$ and $\Delta TT$, which represent the absolute deviation in minutes of the solution’s departure times and travel times, respectively, from the initially requested schedules, provide additional insights into these matters. Smaller deviations in these metrics, as achieved by GA, result in fewer penalties due to schedule changes, making such solutions more favorable in contexts where adherence to initial schedules is critical. However, it is worth noting that larger deviations can sometimes allow for the planning of a greater number of trains, which may lead to a higher revenue for the IM. This trade-off highlights the importance of balancing revenue maximization with adherence to a schedule. In this case, GA excels at minimizing deviations and maximizing revenue compared with the other methods.
@@
Table 6: Summary of results.
Figure 10 shows the statistical comparison based on $p$-values using the Wilcoxon test [63, 64] for paired samples. This test aims to determine whether there are significant differences in the performance between two related samples, making it well suited for comparing algorithm results across paired runs. Values below 0.05 are displayed in green, indicating a significant difference in favor of one algorithm, while values above 0.05 are shown in red, indicating no significant difference. An algorithm is not compared with itself. Therefore, the main diagonal of the matrix has empty values. The results reveal significant differences between most pairs of algorithms at a significance level of 95%. However, certain pairs, such as ACOR versus CMA-ES and Artificial Bee Colony (ABC) exhibit $p$-values greater than 0.05 ($p=0.994$ and $p=0.244$ respectively), indicating no statistically significant difference in their results. Similarly, DE versus WOA yields a $p$-values of $p=0.389$, suggesting comparable performance. These non-significant differences may be due to overlapping performance in certain problem instances, where both methods reach similar solutions.
@@ 4.2 Tests with hyperparameter tuning
In this second test, the two hyperparameters common to all the algorithms used (except SA), the number of iterations and the number of individuals, were adjusted. This was accomplished using the HyperOpt Python package [65, 66].The results from Table 7 illustrate the tuned hyperparameters for each algorithm. These adjusted values were determined based on the exploration range of 100 to 500 epochs and 10 to 100 individuals, with the exception of SA, which is a single-solution algorithm. Instead, the number of iterations for SA was independently set to 45,000 to ensure comparable computational effort. GA and PSO required the maximum number of epochs (500), suggesting their effectiveness relies on extended iterations to achieve optimal results. However, their individual population sizes (70 for GA and 80 for PSO) reflect a balance between population diversity and computational feasibility. Interestingly, the ACOR algorithm uses the smallest population size (10 individuals) but requires the maximum number of epochs (500). This reflects its unique mechanism, where fewer individuals suffice to produce competitive results over prolonged iterations. On the other hand, DE and GWO employ the maximum population size (100 individuals), but with fewer epochs in the case of DE (300).
@@
Table 7: Hyperparameter Tuning Results.
The convergence curves for the second experiment, shown in Figures 12 and 13, reveal significant differences in algorithm performance after hyperparameter tuning compared to the first experiment without tuning. Figure 12 displays the convergence behavior of all 10 algorithms. GA maintains its position as the best-performing algorithm, achieving the highest fitness values across all epochs. However, its convergence is faster and more stable compared to the untuned experiment, demonstrating the impact of optimized hyperparameters. ACOR and DE also show notable improvements, reaching higher fitness values with reduced variability. Figure 13 focuses on the top three algorithms (GA, ACOR, and DE) with tuned hyperparameters. GA continues to dominate with rapid convergence and stabilization at the highest revenue levels. ACOR and DE exhibit competitive performance, with DE converging faster in earlier epochs and ACOR displaying a steady improvement across all iterations. These results contrast with the first experiment, where DE lagged behind PSO in overall performance. The improvements observed in DE further reinforce the benefits of tailored hyperparameters, and also remark that even high-performing algorithms like GA can benefit from careful hyperparameter adjustment to further optimize their performance.
Figure 12: Convergence curves with hyperparameter tuning
Figure 13: Convergence curves of Top 3 algorithms with hyperparameter tuning
@@
Figure 14: Scattered boxplot of population for Top 3 algorithms with hyperparameter tuning
@@
Table 8: Results by run for each of the applied metaheuristics (updated).
@@
Table 9: Summary of results (updated).
Figure 15 illustrates the statistical differences in performance between pairs of algorithms based on their best fitness values across five independent runs. The Kolmogorov-Smirnov test [67, 68] was used to determine whether the distributions of fitness values for each pair of algorithms differ significantly, with a significance level of 95%. The choice of the Kolmogorov-Smirnov test for the second analysis was motivated by the nature of the comparison and the properties of the test itself. Unlike the Wilcoxon signed-rank test, the Kolmogorov-Smirnov test is more suitable for analizing differences between two independent samples, including differences in shape. As with the first experiment, values below 0.05 are displayed in green, indicating a significant difference in favor of one algorithm, while values above 0.05 are shown in red, indicating no significant difference. The results reveal that most algorithm pairs exhibit statistically significant differences ($p$-values $<0.05$). Notably, the GA consistently shows significant superiority over all other algorithms. Similarly, algorithms such as DE, ACOR, and ABC also show significant differences as compared to others.

### v2

@@ 3.3 General scheme of the optimization process
Figure 5: General optimization scheme.
@@ 1. Generating schedule requests:
Let $SS=\{S_{1},\dots,S_{|SS|}\}$ be the set of schedules, where each $S_{i}$ is a schedule. Each service in $S_{i}$ is composed of a set of stops. For each stop, ROBIN indicates the station ($Sta_{i}^{j}$), arrival time ($at_{i}^{j}$) and departure ($dt_{i}^{j}$). Thus, $S_{i}$ can be expressed as:
$S_{i}=\{(Sta_{i}^{1},at_{i}^{1},dt_{i}^{1}),\dots,(Sta_{i}^{|S_{i}|},at_{i}^{|S_{i}|},dt_{i}^{|S_{i}|})\}$ (15)
@@
2. Computing departure time vector: The metaheuristic algorithms will receive a vector with the departure times as input, considering that the original operational times must be complied with (stop time plus travel time). For this purpose, a one-dimension vector containing the departure times of the services in a sorted way is generated using $SS$. This vector is called $DT$ and can be expressed as:
@@
As mentioned above, with the original requested times, it was not feasible to schedule all three services. However, based on the new $ODT$ times this would be feasible, if a five minute minimum safety gap for each service is included. In order to reach this solution, $S_{1}$ has been advanced by half an hour, and the departure of Service 2 has been delayed, also by thirty minutes.
It is important to remark that this might not be the optimal solution: for example, if RU running service one is much more sensitive to penalties compared to the others, it would be better to delay Service 3 and Service 3 while not modifying the requested times of Service 1.
@@
In the following example, based on the previous $ODT$ vector, a definitive schedule is devised. Even with the optimized departure time vector, it is possible to find conflicts between services, making it infeasible to schedule all the requests. For this reason, the heuristic for service planning is also applied. The Marey chart with the final schedule based on the $ODT$ previously shown for the example schedule can be seen in Figure 6. The modified schedule resulting from the $ODT$ vector is set out in Table 3.
Table 3: Final timetable of rail services organized by station.
@@
Figure 6: Mary chart schedule based on $ODT$.
@@ 3.3.1 A heuristic for railway service scheduling
To address the exponential growth caused by checking all possible combinations, the following heuristic technique has been implemented to obtain the vector of trains to be scheduled. The operation of this heuristic is detailed in Algorithm 1. This algorithm aims to schedule a set of services represented by the $ODT$ vector, resolving conflicts between them to produce a final conflict-free scheduled services vector called $B$. Initially, conflicts between the proposed services are identified (Line 3), and those without conflicts are scheduled (Line 4). Then, the benefits associated with the conflicting services are calculated (Line 5). As long as there are services in conflict (Line 6), the algorithm iteratively selects the service with the highest associated benefit (Line 7), schedules it (Line 8), and updates the list of conflicts by removing the newly scheduled services and their corresponding conflicts (Line 9). This approach seeks to maximize total benefits by prioritizing high-value services and minimizing interference.
@@
1:  Input: ODT (Proposed time vector)
2:  Output: B (Scheduled services vector)
3:  C = CONFLICTS_IN(ODT)
4:  B = SCHEDULE_NO_CONFLICTS_IN(C)
5:  R = REVENUES_OF(C)
[…]
7:     S = MAX(R)
8:     B = B + S
9:     C = C - CONFLICTS_BETWEEN(B, S)
10:  end while
@@ 4 Experiments and results
This section presents the results of metaheuristic algorithms for solving the timetabling problem in a liberalized railway market. Section 4.1 analyzes the impact of the selected safety headway and IM bound in the obtained results and provides further insights on the selected values. Section 4.2 shows the results obtained without hyperparameter optimization, and then after optimization (Section 4.3). This enables evaluation of their effectiveness and configuration recommendations. The Python library used is MEALPY Vanthieu2023 , as it is robust and efficient and implements a large number of cutting-edge population-based metaheuristic algorithms. Finally, Section 4.4 shows the results obtained by the metaheristic algorithms in a larger scale instance with 50 service requests, and compares the results with those obtained by classical mathematical approaches such as the SCIP solver.
4.1 Sensitivity analysis between safety margin & IM bounds
In this subsection, a sensitivity analysis will be carried out regarding the safety margin ($\omega$) and IM bound ($\delta$) impact to the revenue performance. Specifically, 2.5, 5, and 10 for safety margins $\delta$, and 30, 45, and 60 minutes for IM bounds $\omega$ has been tested. These combinations allow to observe how increasing the buffer for train spacing and the IM modification margin affect average revenue. Figure 7 presents a heatmap of mean revenue as a function of safety margin (rows) and IM bound (columns). The color gradient clearly indicates that the highest average revenue (darkest blue) occurs when the safety margin is 2.5 minutes and the IM bound is 30 minutes, while the lowest occurs at a 10 minute safety margin with a 60 minute IM bound. In general, revenue tends to decrease as either safety margin or IM bound increases.
Figure 7: Revenue sensitivity to safety margins and infrastructure bounds.
[…]
When examined by individual algorithm, the following patterns emerge:
  • Genetic Algorithm: Safety margin has a highly significant effect ($p<0.0001$), and IM bound is also significant ($p=0.030$). No interaction effect is observed.
  • Particle Swarm Optimization: Safety margin remains significant ($p=0.0015$), whereas IM bound does not ($p=0.293$). Interaction is non-significant.
  • Differential Evolution: Neither safety nor infrastructure factors are significant ($p>0.05$ in all cases).
  • CMA-ES: IM bound is extremely significant ($p<10^{-7}$), while safety is marginal ($p\approx 0.0629$), and no interaction effect is detected.
  • Whale Optimization Algorithm: IM bound shows a significant effect ($p=0.0082$), safety margin does not ($p=0.870$), and there is no interaction.
  • Artificial Bee Colony: Both safety margin ($p=0.007$) and infrastructure bound ($p<10^{-7}$) are significant; interaction is non-significant.
  • Grey Wolf Optimizer: Only IM bound is significant ($p=0.0091$); safety and interaction are not.
  • Ant Colony Optimization (continuous): IM bound is highly significant ($p<10^{-8}$); safety and interaction are not.
  • Simulated Annealing: Neither safety nor infrastructure factors reach significance ($p>0.05$).
  • Hybrid GWO-WOA: Safety margin is significant ($p=0.0199$) and there is a significant safety × infrastructure interaction ($p=0.0146$); IM bound alone is not significant in this hybrid scheme.
In all experiments, a 10 minute safety margin was chosen for operational and safety reasons (allowing sufficient buffer to absorb minor delays without violating headway constraints). The 60 minute modification margin for the IM corresponds to the value that ADIF (the Spanish IM) states as reasonable in its official declaration on network operations Adif24 , ensuring that proposed timetable adjustments remain within practical regulatory limits.
@@ 4.2 Tests without hyperparameter tuning
With the first experiment, all the selected metaheuristics were tested to maximize the IM revenue obtained based on the same 25 schedule requests. In this first test, no hyperparameter optimization was applied, selecting a standard 100-epoch optimization process and 20 individuals (for the population-based metaheuristics) and $100\cdot 20=2000$ epoch for the single-solution-based algorithms, i.e. Simulated Annealing (SA). The default parameters for each algorithm are shown in Table 4.There are no parameters that can be adjusted (apart from epochs and population size) in the implemented MEALPY version of Covariance Matriz Adaptation (CMA-ES), Grey Wolf Optimize (GWO), Whale Optimization Algorithm (WOA) and Hybrid Grey Wolf - Whale Optimization Algorithm (GWO-WOA). Thus, no values are specified in Table 4 for those algorithms.
Table 4: Default parameters in MEALPY for each algorithm.
Algorithm Parameter Default Value Genetic Algorithm (GA) Crossover probability ($pc$) 0.95 Mutation probability ($pm$) 0.025 Selection method Tournament Particle Swarm Optimization (PSO) Local coefficient ($c1$) 2.05 Global coefficient ($c2$) 2.05 Positive constant (alpha) 0.4 Simulated Annealing (SA) Initial temperature 100 Cooling rate 0.99 Differential Evolution (DE) Weighting factor ($wf$) 0.1 Crossover rate ($cr$) 0.9 Ant Colony Optimization Continuous (ACOR) Sample count 25 Intensification factor 0.5 Zeta 1.0 Covariance Matrix Adaptation Evolution Strategy - - Artificial Bee Colony (ABC) Food source limit 25 Grey Wolf Optimizer - - Whale Optimization Algorithm - - Hybrid Grey Wolf - Whale Optimization Algorithm - -
@@
Figure 9: Convergence curves of all algorithms.
Figure 10: Convergence curves of top 3 algorithms.
@@
Figure 11: Scattered boxplot of population.
Table 5 shows the results achieved for each algorithm for each one of the five different runs, ranked by the obtained revenue (higher to lower). The results highlight the clear dominance of GA across all metrics. It consistently achieves the highest revenue over the majority of runs, demonstrating its effectiveness in optimizing the problem at hand. However, an interesting observation arises in runs three and five of GA, where scheduling more trains (nineteen in run five versus eighteen in run three) does not necessarily translate into higher revenue. This indicates that the quality of the solution is not solely dependent on the number of scheduled trains, but also on other factors such as minimizing schedule changes. This happens because, in order to schedule a higher number of services, it is often required to make more significant changes to the requested times, thereby incurring higher penalties. The metrics $\Delta DT$ and $\Delta TT$, which represent the absolute deviation in minutes of the solution’s departure times and travel times, respectively, from the initially requested schedules, provide additional insights into these matters. Smaller deviations in these metrics, as achieved by GA, result in fewer penalties due to schedule changes, making such solutions more favorable in contexts where adherence to initial schedules is critical. However, it is worth noting that larger deviations can sometimes allow for the planning of a greater number of trains, which may lead to a higher revenue for the IM. This trade-off highlights the importance of balancing revenue maximization with adherence to a schedule. In this case, GA excels at minimizing deviations and maximizing revenue compared with the other methods.
Table 5: Results by run for each of the applied metaheuristics.
@@
Figure 12 shows the statistical comparison based on $p$-values using the Wilcoxon test Derrac12 ; Garcia10 for paired samples. This test aims to determine whether there are significant differences in the performance between two related samples, making it well suited for comparing algorithm results across paired runs. Values below 0.05 are displayed in green, indicating a significant difference in favor of one algorithm, while values above 0.05 are shown in red, indicating no significant difference. An algorithm is not compared with itself. Therefore, the main diagonal of the matrix has empty values. The results reveal significant differences between most pairs of algorithms at a significance level of 95%. However, certain pairs, such as ACOR versus CMA-ES and Artificial Bee Colony (ABC) exhibit $p$-values greater than 0.05 ($p=0.994$ and $p=0.244$ respectively), indicating no statistically significant difference in their results. Similarly, DE versus WOA yields a $p$-values of $p=0.389$, suggesting comparable performance. These non-significant differences may be due to overlapping performance in certain problem instances, where both methods reach similar solutions.
@@ 4.3 Tests with hyperparameter tuning
In this second test, the two hyperparameters common to all the algorithms used (except SA), the number of iterations and the number of individuals, were adjusted. This was accomplished using the HyperOpt Python package Bergstra13 ; Hutter19 . The results from Table 7 illustrate the tuned hyperparameters for each algorithm. These adjusted values were determined based on the exploration range of 100 to 500 epochs and 10 to 100 individuals, with the exception of SA, which is a single-solution algorithm. Instead, the number of iterations for SA was independently set to 45000 to ensure comparable computational effort. GA and PSO required the maximum number of epochs (500), suggesting their effectiveness relies on extended iterations to achieve optimal results. However, their individual population sizes (70 for GA and 80 for PSO) reflect a balance between population diversity and computational feasibility. Interestingly, the ACOR algorithm uses the smallest population size (10 individuals) but requires the maximum number of epochs (500). This reflects its unique mechanism, where fewer individuals suffice to produce competitive results over prolonged iterations. On the other hand, DE and GWO employ the maximum population size (100 individuals), but with fewer epochs in the case of DE (300).
Table 7: Hyperparameter tuning results.
@@
The convergence curves for the second experiment, shown in Figures 14 and 15, reveal significant differences in algorithm performance after hyperparameter tuning compared to the first experiment without tuning. Figure 14 displays the convergence behavior of all 10 algorithms. GA maintains its position as the best-performing algorithm, achieving the highest fitness values across all epochs. However, its convergence is faster and more stable compared to the untuned experiment, demonstrating the impact of optimized hyperparameters. ACOR and DE also show notable improvements, reaching higher fitness values with reduced variability. Figure 15 focuses on the top three algorithms (GA, ACOR, and DE) with tuned hyperparameters. GA continues to dominate with rapid convergence and stabilization at the highest revenue levels. ACOR and DE exhibit competitive performance, with DE converging faster in earlier epochs and ACOR displaying a steady improvement across all iterations. These results contrast with the first experiment, where DE lagged behind PSO in overall performance. The improvements observed in DE further reinforce the benefits of tailored hyperparameters, and also remark that even high-performing algorithms like GA can benefit from careful hyperparameter adjustment to further optimize their performance.
Figure 14: Convergence curves of all algorithms with hyperparameter tuning.
Figure 15: Convergence curves of top 3 algorithms with hyperparameter tuning.
@@
Figure 16: Scattered boxplot of population for top 3 algorithms with hyperparameter tuning.
@@
Table 8: Results by run for each of the applied metaheuristics with hyperparameter tunning.
@@
Table 9: Summary of results with hyperparemeter tunning.
@@
Figure 17 illustrates the statistical differences in performance between pairs of algorithms based on their best fitness values across five independent runs. The Kolmogorov-Smirnov test Kolmogorov33 ; Smirnov48 was used to determine whether the distributions of fitness values for each pair of algorithms differ significantly, with a significance level of 95%. The choice of the Kolmogorov-Smirnov test for the second analysis was motivated by the nature of the comparison and the properties of the test itself. Unlike the Wilcoxon signed-rank test, the Kolmogorov-Smirnov test is more suitable for analizing differences between two independent samples, including differences in shape. As with the first experiment, values below 0.05 are displayed in green, indicating a significant difference in favor of one algorithm, while values above 0.05 are shown in red, indicating no significant difference. The results reveal that most algorithm pairs exhibit statistically significant differences ($p$-values $<0.05$). Notably, the GA consistently shows significant superiority over all other algorithms. Similarly, algorithms such as DE, ACOR, and ABC also show significant differences as compared to others.
@@
4.4 Problem scalability
This subsection assesses the scalability of the timetabling problem when addressed by two distinct approaches: metaheuristic algorithms using the MEALPY library and a mathematical programming model implemented in Pyomo and solved with SCIP. As the number of services and stations increases, both the combinatorial complexity of conflict detection and the size of the optimization model grow nonlinearly. Metaheuristic techniques rely on stochastic search mechanisms that can explore and exploit large solution spaces. In contrast, SCIP— currently one of the fastest solvers for Mixed Integer Programming (MIP) and Mixed Integer Nonlinear Programming (MINLP)— employs systematic branching, cutting planes, and presolving procedures to obtain optimal or near‐optimal schedules. By comparing execution times, convergence behavior, and solution quality across problem instances of increasing size, it becomes possible to determine how each methodology scales and to identify practical thresholds for real‐world applications.
A series of benchmark instances is created with 25 and 50 train services on a fixed network topology. For the metaheuristic approach, the Genetic Algorithm (GA) from the mealpy library is set up with an optimized epoch count of 500 and a population size of 70. In contrast, the Pyomo SCIP formulation is solved with a wall‐clock time limit of 2 hours for the 25‐service instance and 8 hours for the 50‐service instance. For each instance, the following metrics are recorded: (a) total wall‐clock time to reach a feasible solution, (b) best revenue achieved within the time limit, and (c) optimality gap relative to SCIP’s best bound when available. This setup permits a direct comparison of GA’s stochastic search performance against SCIP’s exact optimization process as instance size grows.
The metaheuristic results for the 50‐service instance (Table 11) confirm that the Genetic Algorithm (GA) outperforms all other methods by a substantial margin. GA achieves the highest average revenue ($\mu_{\text{revenue}}=6721.32$) and schedules nearly all services on average ($\mu_{\text{trains}}=29$ out of 50), with low variability ($\sigma_{\text{revenue}}=145.21$, $\sigma_{\text{trains}}=1$). It finishes the optimization process (with 500 epochs and 70 individuals) in approximately 571 seconds on average. Differential Evolution (DE) is the second‐best, obtaining $\mu_{\text{revenue}}=5\,215.24$ and $\mu_{\text{trains}}=25$, though with higher spread ($\sigma_{\text{revenue}}=303.03$, $\sigma_{\text{trains}}=2$) in about 483 seconds. Particle Swarm Optimization (PSO) ranks third ($\mu_{\text{revenue}}=4\,921.95$, $\mu_{\text{trains}}=19$) but requires a longer runtime ($\mu_{\text{time}}=643\,\text{s}$) and shows greater variability ($\sigma_{\text{revenue}}=460.84$, $\sigma_{\text{trains}}=2$).
Overall, GA not only finds the highest‐quality schedules but does so with relatively low variance, making it the clear choice among metaheuristics for large‐scale timetable adjustment.
Table 11: Summary of results with 50 requests.
Rank Algorithm $\mu_{time}$ $\sigma_{time}$ $\mu_{revenue}$ $\sigma_{revenue}$ $\mu_{trains}$ $\sigma_{trains}$ 1 Genetic Algorithm 571.01 2.19 6721.32 145.21 29 1 2 Differential Evolution 483.30 1.69 5215.24 303.03 25 2 3 Particle Swarm Optimization 643.06 4.78 4921.95 460.84 19 2 4 Ant Colony Optimization Continuous (ACOR) 220.95 0.87 4555.20 603.54 23 3 5 Covariance Matrix Adaptation Evolution Strategy 961.45 35.75 3997.22 168.33 19 1 6 Whale Optimization Algorithm 505.12 4.71 3890.43 267.76 18 2 7 Grey Wolf Optimizer 776.15 1.81 3794.26 755.06 20 3 8 Artificial Bee Colony 863.28 1.09 3444.28 98.97 20 1 9 Hybrid Grey Wolf - Whale Optimization Algorithm 349.18 0.68 2735.24 70.78 19 2 10 Simulated Annealing 73.06 0.98 2393.08 260.55 16 2
The Pyomo formulation includes nonlinear terms in the objective (penalizing deviations in departure and travel times via Equation 10 as well as bilinear products in the conflict‐avoidance constraints such as Equation 8. This combination of nonlinear penalties and bilinear scheduling variables makes the resulting model a mixed‐integer nonlinear program (MINLP) rather than a pure MIP.
SCIP was chosen because it provides native support for MINLPs, including nonlinear objectives and bilinear constraints, while still remaining non‐commercial. Its branch‐and‐bound engine, advanced presolving techniques, and integrated nonlinear solvers allow it to handle both the smooth penalty functions and the product terms in the conflict constraints. In practice, SCIP was able to produce high‐quality feasible solutions and tighten optimality gaps within the allotted time limits (2 hour for 25 services, 8 hour for 50 services), making it suitable for benchmarking against the GA’s stochastic search.
Table 12 summarizes SCIP’s performance on the 25‐service instance. The original model comprises 83 variables (25 binary, 58 continuous) and 6973 constraints; after presolve, it expands to 383 variables (25 binary, 300 implicits, 58 continuous) and 6565 constraints (including 300 logical “and” and 6265 nonlinear rows). SCIP’s ALNS heuristic finds a nonzero solution ($revenue=445.48$) within 1 second, and subsequent heuristics raise revenue to 4159.14 by 144 seconds, with over 67000 LP iterations. No finite dual bound is obtained before the 600 seconds limit (dual bound reported as $+1\times 10^{20}$), so the optimality gap remains infinite. In practice, the large MINLP structure—dominated by nonlinear conflict and penalty constraints—prevents bound tightening, meaning SCIP produces high‐quality feasible schedules but cannot certify optimality within the time limit.
Table 12: SCIP performance summary on the 25‐service example.
Metric Value
Original variables 83 (25 binary, 58 continuous)
Original constraints 6973
Presolved variables 383 (25 binary, 300 implicits, 58 continuous)
Presolved constraints 6565
Nonlinear constraints 6265
Logical “and” constraints 300
Best primal bound (revenue) 4159.14
Dual bound $+1\times 10^{20}$ (no finite bound)
Optimality gap $\infty$
Total LP iterations 67669
Branch-and-bound nodes processed 8
SCIP status Time limit reached
Wall-clock solve time 7200 (time limit)
For the 50‐service case (Table 13), the model grows to 173 variables (50 binary, 123 continuous) and 26173 constraints, which presolving reduces to 1326 variables (50 binary, 1153 implicits, 123 continuous) and 25024 constraints (including 23 871 nonlinear and 1153 logical “and” rows). SCIP’s heuristics produce a best feasible revenue of 4145.98 but, even after 22625 branch‐and‐bound nodes and over 180 million LP iterations, no finite dual bound is found within the 36000-second limit (dual bound = $+1\times 10^{20}$, gap = $\infty$). The dramatic increase in nonlinear constraints and conflict terms further hampers bound tightening compared to the 25‐service case. Consequently, SCIP can still deliver comparable‐quality feasible schedules, but certifying optimality becomes even less practical as problem size doubles.
Table 13: SCIP performance summary on the 50‐service instance.
Metric Value
Original variables 173 (50 binary, 123 continuous)
Original constraints 26173
Presolved variables 1326 (50 binary, 1 153 implicits, 123 continuous)
Presolved constraints 25024
 Nonlinear constraints 23871
 Logical “and” constraints 1153
Best primal bound (revenue) 4145.98
Dual bound $+1\times 10^{20}$ (no finite bound)
Optimality gap $\infty$
Total LP iterations 180172000
Branch‐and‐bound nodes processed 22625
SCIP status Time limit reached
Wall‐clock solve time 36000 (time limit)
The results in Table 14 highlight clear trade‐offs between GA and Pyomo SCIP. In the 25‐service instance, SCIP proves marginally better revenue (4471.52 vs. 4457.33) but requires over 3000 seconds to reach that solution, whereas GA attains its best in only 174 seconds. Both methods schedule all 19 feasible trains, but GA’s runtime is an order of magnitude lower—showing that metaheuristics can rapidly find near‐optimal schedules.
For the 50‐service instance, GA not only schedules a higher number trains (29) but also achieves significantly higher revenue (6838.61 vs. 4145.98). SCIP, limited by a 36000 seconds time cap, only schedules 21 trains and yields a much lower revenue. This contrast underscores that as problem size increases, SCIP’s inability to close the optimality gap (and its slower convergence) results in suboptimal feasible solutions, whereas GA delivers higher revenue and better coverage in under 600 seconds.
Table 14: Scalability comparison of GA vs. Pyomo SCIP on 25 and 50 service requests.
Metric 25 services 50 services
GA SCIP GA SCIP
Variables Binary: 25 Binary: 50
Continuous: 58 Continuous: 123
Best Revenue 4457.33 4471.52 6838.61 4145.98
Time to Best 174.37s 53min 570.11s 503min
Trains Scheduled 19 19 29 21
Overall, GA shows both computational efficiency and solution quality advantages at larger scales, while SCIP’s exact guarantees are feasible only for smaller instances and come at a substantial time cost.